OpenStudy (anonymous):

After how many visits will Dizneyland and Not's Berry Farm cost the same amount? Not's equation: 20n+60 dizneyland equation: 15n+90 Replace "n" with number of visits.

OpenStudy (danjs):

20n+60=15n+97

OpenStudy (anonymous):

I edited it, sorry, it should've said 90.

OpenStudy (anonymous):

\[20n + 60 = 15n + 90 \] Subtract 60 from both sides \[20n = 15n + 30\] Subtract 15n from both sides \[5n = 30 \] Divide both sides by 5 \[n = 6\]

OpenStudy (dtan5457):

^ beat me to it. that's correct

OpenStudy (anonymous):

Knot's Berry Farm: 20(6) + 60 = 120 + 60 = 180 Disneyland: 15(6) + 90 = 90 + 90 = 180

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem! :)